# Multi-temporal crack segmentation in concrete structure using deep learning approaches

### v1

@@ Abstract
Cracks are among the earliest indicators of deterioration in concrete structures. Early automatic detection of these cracks can significantly extend the lifespan of critical infrastructures, such as bridges, buildings, and tunnels, while simultaneously reducing maintenance costs and facilitating efficient structural health monitoring. This study investigates whether leveraging multi-temporal data for crack segmentation can enhance segmentation quality. Therefore, we compare a Swin UNETR trained on multi-temporal data with a U-Net trained on mono-temporal data to assess the effect of temporal information compared with conventional single-epoch approaches. To this end, a multi-temporal dataset comprising 1356 images, each with 32 sequential crack propagation images, was created. After training the models, experiments were conducted to analyze their generalization ability, temporal consistency, and segmentation quality. The multi-temporal approach consistently outperformed its mono-temporal counterpart, achieving an IoU of $82.72\%$ and a F1-score of $90.54\%$, representing a significant improvement over the mono-temporal model’s IoU of $76.69\%$ and F1-score of $86.18\%$, despite requiring only half of the trainable parameters. The multi-temporal model also displayed a more consistent segmentation quality, with reduced noise and fewer errors. These results suggest that temporal information significantly enhances the performance of segmentation models, offering a promising solution for improved crack detection and the long-term monitoring of concrete structures, even with limited sequential data.
@@ 1 Introduction
Infrastructures such as highways, roads, dams, harbors, and buildings are essential for the sustainability and efficiency of economic activities, ensuring public safety, and facilitating social interactions (Zhou et al.,, 2023). Many of these infrastructures are constructed from materials such as concrete, asphalt, or stone and face various challenges, including fatigue stress, cyclic loading, and the cumulative effects of time, along with increasing human and environmental pressure (Mohan and Poobal,, 2018; König et al.,, 2022; Kheradmandi and Mehranfar,, 2022). These factors can compromise structural integrity (Hamishebahar et al.,, 2022) and diminish the asset value of these constructions (Kheradmandi and Mehranfar,, 2022). Consequently, there is a pressing need for effective Structural Health Monitoring (SHM) to maintain and prolong the lifespan of critical structures (Hamishebahar et al.,, 2022). This helps reduce maintenance and repair costs (Kheradmandi and Mehranfar,, 2022; König et al.,, 2022) and prevents damage to humans and the environment (Mohan and Poobal,, 2018).
[…]
Looking across the Atlantic, a 2022 report highlighted that the United States has $2000$ dams and over 46 0004600046\,00046 000 bridges with structural deficits (König et al.,, 2022), and $11\%$ of roads are classified as being in poor or mediocre conditions (Zhou et al.,, 2023). Similarly, in the United Kingdom, approximately $10\%$ of roads are disrepaired (König et al.,, 2022). These examples illustrate the significant global need for infrastructure assessment and maintenance.
One of the earliest and most common indicators of structural degradation and reduced structural integrity is the presence of cracks on the surfaces of structures (König et al.,, 2022; Hamishebahar et al.,, 2022). Cracks begin at the microscopic level and continuously reduce the local stiffness of materials and create material discontinuities (Mohan and Poobal,, 2018). If left untreated, these cracks grow in size, and the cost and effort required to repair them increase accordingly. To mitigate consequential safety issues, it is crucial for SHM to precisely assess the state of a structure and identify indicators of future damage such as cracks König et al., (2022).
@@
To the best of the authors’ knowledge, as stated by König et al. (2022), there is a lack of research regarding the use of multi-temporal data for CV tasks related to crack detection. In this context, we focus on the semantic segmentation of cracks using multi-temporal data, which is compared to approaches that utilize mono-temporal data. In that regard, a multi-temporal dataset and a corresponding deserialized mono-temporal dataset were created. Moreover, we aim to address the research gap in multi-temporal crack propagation data for semantic segmentation. Additionally, the development of a multi-temporal dataset can provide a foundation for future research involving multi-temporal data. The contributions of this study can be summarized as follows:
  • Development of mono- and multi-temporal models for crack segmentation.
  • Assessment of mono- and multi-temporal models based on performance metrics.
@@ 2 Related works
Crack detection techniques can be based on RGB, infrared, ultrasonic, laser, and other types of images (Mohan and Poobal,, 2018). In this study, we focused on camera-based detection, similar to the dataset used in our experiments. Many review papers categorize automatic crack detection methods into two fundamental types: rule-based and data-driven approaches (Kheradmandi and Mehranfar,, 2022; Gong et al.,, 2024; Hamishebahar et al.,, 2022). Rule-based methods rely on edge information of cracks, morphological operations, and thresholding techniques. These methods may struggle to perform adequately on less structured and more complex data because of their rigidity and inability to adapt to different contexts. By contrast, data-driven approaches outperform rule-based approaches in managing complex scenes by learning useful feature representations from training data, particularly DL-based methods. As a result, they tend to be more accurate, easier to automate (Gong et al.,, 2024) and computationally more efficient. Consequently, this study focused on exploring DL approaches.
@@
Çiçek et al. (2016) introduced a 3D U-Net, where 3D convolutional layers are employed to segment volumetric images. This model can be easily extended to multi-temporal data, where the third dimension is interpreted not as depth, but as the temporal domain. Xu et al. (2019) presented a Long Short-Term Memory (LSTM) multi-modal U-Net, which is a U-Net with multiple encoder paths designed to process multi-modal data for the segmentation of brain tumor data. An attempt was made to increase the learning capabilities of the model by not fusing the feature extraction of different channels. The decoder concatenates information from every encoder path and produces a segmentation mask. This multi-modal U-Net is applied to different slices of volumetric multi-modal input data. The output segmentation masks create a sequence of segmentation masks based on the depth of the image. This sequence is sequentially input into Convolutional LSTM (ConvLSTM) networks to capture the sequence dependencies. Therefore, volumetric data is fully segmented. Rustowicz et al. (2019) employed a 2D U-Net in combination with a ConvLSTM for crop type segmentation using multi-modal and multi-temporal satellite data. The model has a U-Net type structure, where the encoder first extracts the features of each channel independently using CNNs, and these features are each fed into a ConvLSTM. Zhou et al. (2023) proposed the Swin UNETR model, which also employs a U-Net architecture with a Swin Transformer as a backbone in the encoder, enhanced to process volumetric data. Feature extraction of the backbone relies solely on the attention mechanism, whereas CNNs are only employed within residual blocks in skip connections to refine the features. The decoder concatenates features from the feature maps with different resolutions and upsamples them to the size of the input data. Therefore, volumetric or multi-temporal data can be fully segmented. Because this model extracts features using the attention mechanism and can segment multi-temporal data in the original input size, we selected this model for multi-temporal crack segmentation. Moreover, to the best of the authors’ knowledge, attention mechanisms have not yet been applied to multi-temporal crack segmentation.
@@ 3.1 Dataset
[…]
@@
Figure 3: Swin UNETR architecture. A 3D Swin Transformer is used as a feature extractor, and in the decoder, the feature maps are concatenated and upsampled to the original input size. Modified from Hatamizadeh et al. (2022).
The creation of mono- and multi-temporal datasets can be divided into three phases: labeling, post-processing, and aggregation. In the labeling phase, because manual annotation of all images is tedious and extremely time-consuming, a set of three images was manually annotated and used to train a segmentation model (auxiliary U-Net) to assist the annotation process of the remaining 25 images. The post-processing phase involves the removal of false positives (pencil drawings, concrete pores, shadows, and cables) and segmentation errors (isolated groups of pixels smaller than five pixels) in the auxiliary U-Net predictions. Then, morphological operations were applied (closing with a $3\times 3$ kernel), followed by manual cleaning to refine the crack annotations. Finally, the aggregation phase consists of fulfilling the model requirements for training, as the multi-temporal model requires sequences of 32 images, but only 25 are available. To address this, seven images at evenly spaced intervals were duplicated to expand the sequence to 32 images.
To create the mono-temporal dataset, patches of $128\times 128$ were extracted from the image sequences and their corresponding annotations (targets), resulting in a total of 180 224 image-target pairs. The multi-temporal dataset contains a stack of all images, resulting in 5 632 multi-temporal image-target pairs. As expected, there were several samples without cracks. To balance the dataset, all multi-temporal samples with cracks were included, and samples without cracks were randomly selected to achieve a 2:1 ratio of images with cracks to those without cracks. The training, validation, and test sets were created at the following proportions: 60%, 20%, and 20%, respectively. To generate the mono-temporal dataset and ensure that both models have access to exactly the same data, the multi-temporal samples within the splits were deserialized. Tables 1 and 2 show the dataset statistics.
Sample size Crack image ratio (%) Crack pixel ratio (%)
All 43 3924339243\,39243 392 $40.1$ $1.2$
Train 26 0162601626\,01626 016 $40.1$ $1.2$
Val 8 67286728\,6728 672 $39.1$ $1.3$
Test 8 70487048\,7048 704 $41.3$ $1.2$
Table 1: Mono-temporal dataset. The deserialization of the multi-temporal dataset leads to a decreased crack image ratio in the mono-temporal dataset.
Sample size Crack image ratio (%) Crack pixel ratio (%)
All 1 35613561\,3561 356 $66.7$ $1.2$
@@
Table 2: Multi-temporal dataset. The sample size of the multi-temporal dataset is $32$ times smaller than in the mono-temporal dataset because of sequential stacking.
@@ 3.2 Methodology
Swin UNETR: The Shifting Window U-Net Transformer (Swin UNETR) was introduced by Hatamizadeh et al. (2022) and extends the Shifting Window Transformer (Swin Transformer) (Liu et al.,, 2021) to perform semantic segmentation on three-dimensional inputs. This is achieved by combining a Swin Transformer adapted for three-dimensional inputs with a U-Net architecture. In Figure 3 the architecture of Swin UNETR is depicted. The input tokens are patches of pixels, which are embedded into a feature dimension $C$ in the input layer. The Swin Transformer then serves as a feature extractor in the encoder path, whereas the decoder path reassembles the original input shape and predicts each voxel using the feature maps produced by the Swin Transformer. The Swin Transformer enables dense prediction tasks for the transformer architecture, which is not feasible in the preceding Vision Transformer (ViT) (Dosovitskiy et al.,, 2021), owing to its quadratic computational cost.
@@ 4.1 Experimental Setup
As baselines, the following models were used: a mono-temporal U-Net (31M parameters) and a multi-temporal SwinUNETR (15.7M parameters), both implemented using PyTorch (Paszke et al.,, 2019) and trained on a Tesla P100 GPU. We employed early stopping with a patience of 20 epochs, monitoring the validation Intersection over Union (IoU) for the crack class, along with a step learning rate scheduler that adjusted the learning rate by a factor of $0.1$ if the validation loss for the class crack does not decrease for $10$ epochs. For the batch size, a value of $4$ was chosen. Both models processed $128\times 128$ images owing to memory constraints. SwinUNETR was taken from the Medical Open Network for AI (MONAI) (Cardoso et al.,, 2022) framework and has an input channel size of $3$, an output channel size of $1$, and a feature size $C=24$; otherwise, the configuration is the default one described in the MONAI documentation.
@@ 4.2.1 Quantitative analysis:
BL Swin UNETR $\mathbf{82.7}$ $\mathbf{91.9}$ $89.2$ $\mathbf{90.5}$
$1$ U-Net $71.9$ $80.5$ $87.1$ $83.7$
$2$ Swin UNETR $82.5$ $91.2$ $\mathbf{89.6}$ $90.4$
$3$ U-Net $75.7$ $86.0$ $86.3$ $86.2$
@@ 4.2.2 Qualitative analysis:
Figure 5: Early stage vs late stage crack segmentation. Both models perform well on early stages cracks, but only the Swin UNETR manages to maintain adequate segmentation quality in later stage cracks.
U-Net and Swin UNETR demonstrated similar performance on early stage cracks, as illustrated in Figure 5, where both models produce consistent predictions that accurately depict the target. However, in the later stages of crack propagation, as shown in the subsequent images of Figure 5, U-Net’s predictions become less consistent and exhibit numerous holes not present in the target. In contrast, the Swin UNETR maintains a high level of consistency with no holes in its segmentation.
4.2.3 Evaluation of false positives:
On the concrete slab, there were several distinct challenges for both models to handle, such as pencil-drawn numbers, sensors, cables, thin and low-contrast cracks, and cavities which all proved to be potential sources for false positive classifications. By generating the prediction of the entire image of the concrete slab, we can see whether the models learned to mitigate these challenges.

### v2

# Multi-temporal crack segmentation in concrete structures using deep learning approaches

@@ Abstract
Cracks are among the earliest indicators of deterioration in concrete structures. Early automatic detection of these cracks can significantly extend the lifespan of critical infrastructures, such as bridges, buildings, and tunnels, while simultaneously reducing maintenance costs and facilitating efficient structural health monitoring. This study investigates whether leveraging multi-temporal data for crack segmentation can enhance segmentation quality. Therefore, we compare a Swin UNETR trained on multi-temporal data with a U-Net trained on mono-temporal data to assess the effect of temporal information compared with conventional single-epoch approaches. To this end, a multi-temporal dataset comprising 1356 images, each with 32 sequential crack propagation images, was created. After training the models, experiments were conducted to analyze their generalization ability, temporal consistency, and segmentation quality. The multi-temporal approach consistently outperformed its mono-temporal counterpart, achieving an IoU of $82.72\%$ and a F1-score of $90.54\%$, representing a significant improvement over the mono-temporal model’s IoU of $76.69\%$ and F1-score of $86.18\%$, despite requiring only half of the trainable parameters. The multi-temporal model also displayed a more consistent segmentation quality, with reduced noise and fewer errors. These results suggest that temporal information significantly improves the performance of segmentation models, offering a promising solution for improved crack identification and long-term monitoring of concrete structures, even with limited sequential data. The dataset is available at: https://github.com/saidharb/Multi-Temporal-Crack-Segmentation-Dataset.git
@@ 1 Introduction
Infrastructures such as roads, bridges, dams, harbors, and buildings are essential to the sustainability and efficiency of economic activities, to ensure public safety, and to facilitate social interactions (Zhou et al.,, 2023; Maboudi et al.,, 2024). Many of these infrastructures are constructed from materials such as concrete, asphalt, or stone and face various challenges, including fatigue stress, cyclic loading, and cumulative effects of time, along with increasing human and environmental pressure (Mohan and Poobal,, 2018; König et al.,, 2022; Kheradmandi and Mehranfar,, 2022). These factors can compromise structural integrity (Hamishebahar et al.,, 2022) and diminish the asset value of these constructions (Kheradmandi and Mehranfar,, 2022). Consequently, there is a pressing need for effective Structural Health Monitoring (SHM) to maintain and prolong the lifespan of critical structures (Hamishebahar et al.,, 2022). This helps reduce maintenance and repair costs (Kheradmandi and Mehranfar,, 2022; König et al.,, 2022) and prevents damage to humans and the environment (Mohan and Poobal,, 2018).
Numerous structures worldwide are in poor condition and require continuous monitoring and damage detection systems. For instance, German seaports feature approximately $3000\,\text{km}$ quay walls and $2500\,\text{km}$ facilities along the federal waterways. Throughout $140\,\text{km}$ of these quay walls, approximately $279$ million tons of goods are handled, underscoring the significant role of these structures in the German economy. However, the current state of quay walls and facilities is a concern. According to the Federal Waterways Engineering and Research Institute (BAW), $70\%$ of quay walls and facilities on waterways are in adequate condition. This situation foreshadows the considerable maintenance efforts required to ensure continued transportation of essential goods on German waterways (Alamouri et al.,, 2024).
Looking across the Atlantic, a 2022 report highlighted that the United States has $2000$ dams and over 46 00046\,000 bridges with structural deficits (König et al.,, 2022), and $11\%$ of roads are classified as in poor or mediocre conditions (Zhou et al.,, 2023). Similarly, in the United Kingdom, approximately $10\%$ of roads are disrepaired (König et al.,, 2022). These examples illustrate the significant global need for infrastructure assessment and maintenance.
One of the first and most common indicators of structural degradation and reduced structural integrity is the presence of cracks on the surfaces of structures (König et al.,, 2022; Hamishebahar et al.,, 2022). Cracks begin at the microscopic level and continuously reduce the local stiffness of materials and create material discontinuities (Mohan and Poobal,, 2018). If left untreated, these cracks grow in size, and the cost and effort required to repair them increase accordingly. To mitigate consequential safety issues, it is crucial for SHM to accurately assess the state of a structure and identify indicators of future damage, such as cracks König et al., (2022).
@@
To the best of the authors’ knowledge, and as stated by König et al. (2022), there is a lack of research regarding the use of multi-temporal data for CV tasks related to crack detection. In this context, we focus on the semantic segmentation of cracks using multi-temporal data, which is compared to approaches that utilize mono-temporal data. We aim to address the research gap in multi-temporal crack propagation data for semantic segmentation by providing a multi-temporal dataset and a corresponding deserialized mono-temporal dataset, which provide a foundation for future research. The contributions of this study can be summarized as follows:
  • Development and assessment of mono- and multi-temporal models for crack segmentation.
  • Creation of a multi-temporal dataset and a corresponding deserialized mono-temporal dataset.
@@ 2 Related works
Crack detection techniques can be based on RGB, infrared, ultrasonic, laser, and other types of images (Mohan and Poobal,, 2018). In this study, we focused on camera-based detection, similar to the dataset used in our experiments. Many review papers categorize automatic crack detection methods into two fundamental types: rule- and data-driven approaches. Rule-driven methods rely on edge information of cracks, morphological operations, and thresholding techniques. These methods may struggle to perform adequately on less structured and more complex data because of their rigidity and inability to adapt to different contexts (Kheradmandi and Mehranfar,, 2022; Gong et al.,, 2024; Hamishebahar et al.,, 2022). By contrast, data-driven approaches outperform rule-driven approaches in managing complex scenes by learning useful feature representations from training data, particularly DL-based methods. As a result, they tend to be more accurate, easier to automate (Gong et al.,, 2024) and computationally more efficient. Consequently, this study focused on exploring DL approaches.
@@
Çiçek et al. (2016) introduced a 3D U-Net, where 3D convolutional layers are employed to segment volumetric images. This model can be easily extended to multi-temporal data, where the third dimension is interpreted not as depth, but as the temporal domain. Xu et al. (2019) presented a Long Short-Term Memory (LSTM) multi-modal U-Net, which is a U-Net with multiple encoder paths designed to process multi-modal data for the segmentation of brain tumor data. An attempt was made to increase the learning capabilities of the model by not fusing the feature extraction of different channels. The decoder concatenates information from every encoder path and produces a segmentation mask. This multi-modal U-Net is applied to different slices of volumetric multi-modal input data. The output segmentation masks create a sequence of segmentation masks based on the depth of the image. This sequence is sequentially input into Convolutional LSTM (ConvLSTM) networks to capture the sequence dependencies. Therefore, volumetric data is fully segmented. Rustowicz et al. (2019) employed a 2D U-Net in combination with a ConvLSTM for crop type segmentation using multi-modal and multi-temporal satellite data. The model has a U-Net type structure, where the encoder first extracts the features of each channel independently using CNNs, and these features are each fed into a ConvLSTM. Zhou et al. (2023) proposed the Swin UNETR model, which also employs a U-Net architecture with a Swin Transformer as a backbone in the encoder, enhanced to process volumetric data. Feature extraction of the backbone relies solely on the attention mechanism, whereas CNNs are only employed within residual blocks in skip connections to refine the features. The decoder concatenates features from the feature maps with different resolutions and upsamples them to the size of the input data. Therefore, volumetric or multi-temporal data can be fully segmented. We selected this model for multi-temporal crack segmentation because it uses attention mechanisms to extract features and segment data at the original input size. To the best of the authors’ knowledge, attention mechanisms have not yet been applied to multi-temporal crack segmentation.
@@ 3.1 Dataset
A temporal sequence of images depicting crack propagation is required to create a multi-temporal dataset for crack segmentation. For this purpose, a setup was constructed involving a vertical concrete slab with an area of $2\,\text{m}^{2}$ and thickness of $15\,\text{cm}$ (see Figure 2), which was reinforced and positioned upright. A force was applied to the center of the slab, with the load incrementally increasing by $20\,\text{kN}$ at each load stage, referred to as epochs, until the maximum load was reached. Eight epochs were recorded before the concrete slab failed. During each epoch, high-resolution RGB images, aligned by reference points on the concrete slab, were captured using a mounted camera positioned at a specified distance (Backhaus et al.,, 2024). This study used $25$ images with a resolution of $11664\times 8750$ pixels and a ground sampling distance of $0.3\,\text{mm}$ from the final epoch to generate training data. These images are in temporal order and provide a detailed view of the entire concrete block, with the progression of cracks between the images visible upon visual inspection.
@@
Figure 3: Swin UNETR architecture. A 3D Swin Transformer is used as a feature extractor, and in the decoder, the feature maps are concatenated and upsampled to the original input size. Adopted from (Hatamizadeh et al.,, 2022).
The creation of mono- and multi-temporal datasets can be divided into three phases: labeling, post-processing, and aggregation. In the labeling phase, because manual annotation of all images is tedious and extremely time-consuming, a set of three images was manually annotated and used to train a segmentation model (auxiliary U-Net) to assist the annotation process of the remaining 25 images. The post-processing phase involves the removal of false positives (pencil drawings, concrete pores, shadows, and cables) and segmentation errors (isolated groups of pixels smaller than five pixels) in the auxiliary U-Net predictions. Then, morphological operations were applied (closing with a $3\times 3$ kernel), followed by manual cleaning to refine the crack annotations. Finally, the aggregation phase consists of fulfilling the model requirements for training, as the multi-temporal model requires sequences of 32 images (cf. Section 3.2), but only 25 are available. To address this, seven images at evenly spaced intervals were duplicated to expand the sequence to 32 images.
Type
Sample
size
Crack image
ratio (%)
Crack pixel
ratio (%)
Multi-
temp.
All 1 3561\,356 $66.7$ $1.2$
@@
Mono-
temp.
All 43 39243\,392 $40.1$ $1.2$
Train 26 01626\,016 $40.1$ $1.2$
Val 8 6728\,672 $39.1$ $1.3$
Test 8 7048\,704 $41.3$ $1.2$
Table 1: Statistics for multi-temporal and mono-temporal datasets.
To create the multi-temporal dataset, images and corresponding segmentation targets were stacked in temporal order. From these stacks, non-overlapping spatial patches of size $128\times 128$ pixels were extracted, resulting in a total of 5,632 samples, each consisting of a sequence of 32 consecutive crack propagation image-target pairs. As expected, there were several samples without cracks. To balance the dataset, all multi-temporal samples with cracks were included, and samples without cracks were randomly selected and removed to achieve a 2:1 ratio of images with cracks to those without cracks. The training, validation, and test sets were created at the following proportions: 60%, 20%, and 20%, respectively. To generate the mono-temporal dataset and ensure that both models have access to exactly the same data, the multi-temporal samples within the splits were deserialized. Table 1 shows the dataset statistics.
The sample size of the multi-temporal dataset is $32$ times smaller than in the mono-temporal dataset because of sequential stacking. The deserialization of the multi-temporal dataset leads to a decreased crack image ratio in the mono-temporal dataset.
@@ 3.2 Methodology
Swin UNETR: The Shifting Window U-Net Transformer (Swin UNETR) was introduced by Hatamizadeh et al. (2022) and extends the Shifting Window Transformer (Swin Transformer) (Liu et al.,, 2021) to perform semantic segmentation on three-dimensional inputs. This is achieved by combining a Swin Transformer adapted for three-dimensional inputs with a U-Net architecture. In Figure 3 the architecture of Swin UNETR is depicted. The input tokens are patches of pixels, which are embedded into a feature dimension $C$ in the input layer. The Swin Transformer then serves as a feature extractor in the encoder path, whereas the decoder path reassembles the original input shape and predicts each voxel using the feature maps produced by the Swin Transformer. The encoder path features five downsampling operations, resulting in a reduction of resolution by a factor of $2^{5}=32$. Therefore the minimum input resolution of samples is $32\times 32\times 32$ pixels. The Swin Transformer enables dense prediction tasks for the transformer architecture, which is not feasible in the preceding Vision Transformer (ViT) (Dosovitskiy et al.,, 2021), owing to its quadratic computational cost.
@@
U-Net: The U-Net was introduced by Ronneberger et al. (Ronneberger et al.,, 2015) and features a encoder-decoder structure employing convolutional down- and up-sampling layers as well as skip connections. Due to its simple architecture and robust performance it is often used in crack segmentation (Hamishebahar et al.,, 2022; Kheradmandi and Mehranfar,, 2022) and will be considered as a baseline. To produce outputs of the same size as the inputs, padding was aplied in this work.
@@ 4.1 Experimental Setup
As baselines, the following models were used: a mono-temporal U-Net (31M parameters) and a multi-temporal SwinUNETR (15.7M parameters), both implemented using PyTorch (Paszke et al.,, 2019) and trained on a Tesla P100 GPU. We employed early stopping with a patience of 20 epochs, monitoring the validation Intersection over Union (IoU) for the crack class, along with a step learning rate scheduler that adjusted the learning rate by a factor of $0.1$ if the validation loss for the class crack does not decrease for $10$ epochs. For the batch size, a value of $4$ was chosen. Both models processed $128\times 128$ images owing to memory constraints. SwinUNETR was taken from the Medical Open Network for AI (MONAI) (Cardoso and et al.,, 2022) framework and has an input channel size of $3$, an output channel size of $1$, and a feature size $C=24$; otherwise, the configuration is the default one described in the MONAI documentation.
@@ 4.2.1 Quantitative analysis:
BL
Swin UNETR
$\mathbf{82.7}$ $\mathbf{91.9}$ $89.2$ $\mathbf{90.5}$
11 U-Net $71.9$ $80.5$ $87.1$ $83.7$
22
Swin UNETR
$82.5$ $91.2$ $\mathbf{89.6}$ $90.4$
33 U-Net $75.7$ $86.0$ $86.3$ $86.2$
@@ 4.2.2 Qualitative analysis:
Figure 5: Early stage vs. late stage crack segmentation. Both models perform well on early stages cracks except for thin and low contrast cracks, which are also seldom in the targets. Only the Swin UNETR manages to maintain adequate segmentation quality in later stages of crack propagation.
U-Net and Swin UNETR demonstrated similar performance on early stage cracks, as illustrated in Figure 5, where both models produce consistent predictions that accurately depict the target. However, thin and low contrast cracks are a challenge for both models. In the later stages of crack propagation, as shown in the subsequent images of Figure 5, U-Net’s predictions become less consistent and exhibit numerous holes not present in the target. In contrast, the Swin UNETR maintains a high level of consistency with no holes in its segmentation.
4.2.3 Evaluation of false positives and false negatives:
On the concrete slab, there were several distinct challenges for both models to handle, such as pencil-drawn numbers, sensors, cables, thin and low-contrast cracks, and cavities which all proved to be potential sources for false positive classifications. Generating predictions for the entire concrete slab reveals whether the models learned to handle these challenges.